x-3y+7z+w=11
2x+4y+6z-3w=-3
3x+2y+z+2w=19
4x+y-3z+w=22
Systems of linear equations in three variables

Answer by Alan3354(69443)

x-3y+7z+w=11
2x+4y+6z-3w=-3
3x+2y+z+2w=19
4x+y-3z+w=22
----------------
Looks like 4.
Determinant = -535
w = 3
x = 5
y = -1
z = 0
-------
I made an Excel sheet to do these. Doing them manually is tedious and error-prone.
There are sites that will solve up to a 10 by 10 that I've seen.